OpenStudy (bbb911):

If Tucson's average rainfall is 11 1/4 inches and Yuma's is 3 3/5 inches, how much more rain, on the average, does Tucson get then Yuma?

OpenStudy (anonymous):

\[11\frac{ 1 }{ 4 }=\frac{ 44+1 }{ 4 }=\frac{ 45 }{ 4 }\] nd \[3\frac{ 3 }{ 5 }=\frac{ 15+3 }{ 5}=\frac{ 18 }{ 5 }\] just subtract that \[\frac{ 45 }{ 4 }-\frac{ 18 }{ 5 }=\frac{ 225 -72}{ 20 }=\frac{ 153 }{ 20 }\]

OpenStudy (bbb911):

it was close the denominator is right but the numerator isn't

OpenStudy (anonymous):

u gotta break it

OpenStudy (anonymous):

\[7\frac{ 13 }{ 20 }\]

OpenStudy (bbb911):

oh u mean if i divide 153 to 20 right?

OpenStudy (bbb911):

thank you so much :)

OpenStudy (anonymous):

ur welcme ;)